OpenStudy (anonymous):

What is the potential energy stored in a spring that has a restoring force of 1kN/m and has been stretched by 1 cm? Hints: Potential Energy =FS (force x distance stretched) A. 1J B. 10J C. 100J D. 1000J E. No correct answer is available

OpenStudy (anonymous):

You're teacher's hint is faulty. Let me explain. From the Work-Energy Theorem, we know that\[W = \Delta E = \Delta KE + \Delta PE\]This states that work equals the change in energy. In this case, we can say that \(\Delta KE = 0\). Additionally, we know that the definition of work is\[W = \int\limits F \cdot dx\]The force of a spring is\[F_s = kx\]\[\therefore W = \int\limits kx dx\]Integrating obtains\[W = {1 \over 2} kx^2\] Therefore, we can express the potential energy of the spring as\[W =\Delta PE \rightarrow PE_s = {1 \over 2} kx^2\]where k is the spring constant (units of N/m) and x is the displacement (units of m). This will give you units of Joules.